# Prediction of the Most Fire-Sensitive Point in Building Structures with Differentiable Agents for Thermal Simulators

### v1

@@ VII-B2 Evaluation results
TABLE VII: Key CDF values (%) of different metrics to evaluate the MFSP Predictor (numbers in boldface indicate best results).
@@
Comparing the De novo training and TL with GNN methods, we observe that the TL with GNN approach outperforms De novo training in all metrics. For both small and large predictors, TL with GNN achieves lower average distance error, average room distance error, and average room rank, while also attaining higher average MIDR values compared to De novo training using the same loss function. This indicates that the TL with GNN approach effectively leverages the information embedded in the MIDR Predictor’s GNN module, significantly enhancing the performance of the MFSP Predictor.
@@
We compare the MSE loss with Hybrid loss within the TL with GNN method, using the Hybrid ($w_{1}=50$) case as an example. For the average MIDR metric, the Hybrid loss explicitly calculates MIDR, achieving a higher value than the MSE loss by up to $(2.46-2.40)\left/2.40\right.=2.50\%$ for the small predictor, with a smaller improvement observed for the large predictor. Conversely, as the MSE loss is computed based on the pseudo ground truth MFSP (similar to the distance error metric), it achieves a lower average distance error than the that of the Hybrid loss by up to $16\%$ for the large predictor. Interestingly, despite the average room distance error and room rank being calculated relative to the pseudo ground truth MFSP, the Hybrid loss performs similarly to or even better than the MSE loss. This can likely be attributed to the pseudo labels being accurate at the room-level granularity, even though the MFSP is not necessarily located at the exact center of a room. As a result, the MSE and Hybrid loss methods show varying performances across the distance error and room distance error metrics.
@@
The Hybrid loss introduces a trade-off between achieving higher MIDR and lower distance error by adjusting the weights $w_{1}$ and $w_{2}$, which reflect the degree of reliance on the pseudo labels. In this analysis, the weight for the MSE term is fixed at $w_{2}=1$. By varying the weight for the MIDR term ($w_{1}$) from 10 to 100, the reliance on pseudo labels decreases as $w_{1}$ increases. This trade-off is evident when comparing results for $w_{1}=10$ and $w_{1}=50$: $w_{1}=10$ results in approximately $10\%$ lower distance error but also yields a slightly lower MIDR, reduced by about $1\%$. However, when $w_{1}$ is set to a very large value, such as $w_{1}=100$, the reliance on the pseudo labels becomes minimal, and the difficulty of training an “argmaxer”, as discussed in Section VI-A, negatively impacts the MFSP Predictor’s performance. Notably, even for the average MIDR metric, the performance with $w_{1}=100$ is worse than that achieved with $w_{1}=50$.
@@
Taking TL with GNN and Hybrid loss ($w_{1}=50$) as an example, the average room distance error reaches 1.15, indicating that the predicted MFSP is, on average, located within 1.15 rooms of the ground truth MFSP. Additionally, the average room rank is $6.73$ indicating that the predicted MFSP is approximately the 6th or 7th most fire-sensitive room among all rooms in the building structure, with an average of over 87 rooms in each structure in this study. These low values suggest that the MFSP Predictor can accurately identifies the MFSPs in the building.
@@
To further illustrate the performance of the MFSP Predictor, we plotted the Cumulative Distribution Function (CDF), defined as ${\mathcal{F}}_{X}(a)={\mathcal{P}}(X\leq a)$, where $\mathcal{P}$ indicates the probability and $X$ is a random variable representing the different metrics, i.e., distance error, room distance error, and room rank. Figures 10(a to c) display the results for the small predictor, while Figures 10(d to f) present the results for the large predictor. The key CDF values are also summarized in Table VII. In Figure 10, the CDF curves of TL with GNN generally lie above those of De novo training, demonstrating that the TL with GNN approach outperforms De novo training in most cases. Besides, TL with GNN requires much less computational resources in training to achieve this advantage. This is because it leverages a pre-trained GNN, which significantly reduces the computational burden of training from scratch. The pre-trained model provides a well-initialized starting point, requiring only fine-tuning to adapt to the target task. This further highlights the efficiency and superiority of the approach of TL with GNN.
@@
Specifically, as detailed in Table VII, for the small network using TL with GNN and Hybird loss ($w_{1}=50$), $65.40\%$ of cases achieve a distance error of less than $5$ meters. When the threshold is extended to $10$ meters, this ratio increases to $83.21\%$, surpassing the $82.19\%$ achieved with De novo training. Regarding room distance error, $\tilde{e}_{m}$, the large network using TL with GNN and Hybird loss achieves the highest probability of $73.93\%$ for $\tilde{e}_{m}\leq\sqrt{2}$, indicating that the predicted MFSP is either within the same room as the pseudo ground truth MFSP or an adjacent room. Notably, for normalized rooms treated as unit-length squares (Section VII-B1), the longest distance between the centers of two diagonally adjacent rooms is $\sqrt{2}$. When the error threshold is further relaxed to $\tilde{e}_{m}\leq 2$, this probability increases to $80.32\%$, reflecting a high level of accuracy. For room rank, the large network employing TL with GNN and Hybird loss achieves the highest Top-5 accuracy, defined as the fraction of instances where the predicted MFSR ranks among the top 5 pseudo ground truth MFSRs, i.e., ${\mathcal{P}}\left(r_{m}\leq 5\right)$. This accuracy reaches $74.06\%$, representing an improvement of $2.49\%$ over De novo training. Additionally, when using MSE loss, the TL with GNN method achieves a Top-5 accuracy of $72.28\%$, which is $3.7\%$ higher than that of De novo training. These results highlight the robustness of the TL with GNN approach, particularly when paired with Hybrid loss, in accurately identifying the most fire-sensitive locations within building structures.
@@
In summary, the TL with GNN approach consistently outperforms De novo training across all evaluation metrics, highlighting the effectiveness of leveraging the MIDR Predictor’s GNN module to enhance the performance of the MFSP Predictor. Two notable trade-offs arise when selecting the loss function. The first is between $w_{1}$ and $w_{2}$ within the Hybrid loss to include reliance on the room-level granularity of pseudo ground truth MFSP. The second is between MSE loss and Hybrid loss for higher efficiency and accuracy. While the Hybrid loss achieves slightly better results in most metrics (except for distance error), it incurs a higher computational cost due to the additional GNN forward pass required to calculate the MIDR loss and the backward pass through MIDR Predictor to calculate the gradient of the loss $L_{\text{MIDR}}$. In contrast, the MSE Loss avoids this additional computational cost, making it more efficient but potentially less accurate in certain scenarios. This trade-off between computational cost and accuracy should be carefully considered based on the specific requirements and constraints of the application. If computational efficiency is a priority, the MSE loss may be more suitable. However, if higher accuracy in identifying sensitive fire points is critical, the Hybrid loss could provide a more robust solution.
VIII Concluding Remarks
@@
This study introduces a neural network-based framework for predicting the Most Fire-Sensitive Point (MFSP) in building structures, advancing fire safety assessment through a novel combination of data-driven and simulation-based approaches. By employing a differentiable agent model, the proposed framework efficiently simulates Finite Element Analysis (FEA) to predict the Maximum Interstory Drift Ratio (MIDR) under fire conditions. This agent model facilitates rapid generation of MFSP-labeled data, directly supporting the training of the MFSP Predictor. Complementing this, a rule-based temperature generation method was developed, enabling the creation of extensive structural and fire simulation datasets. These contributions establish a comprehensive workflow for data generation and neural network training.
Key innovations include the application of Graph Neural Networks (GNNs) to process structural data, the integration of Transfer Learning (TL) to enhance prediction performance, and the introduction of an Edge Update (EU) mechanism to account for property changes in beams and columns during fire scenarios. Experimental evaluations demonstrated the framework’s effectiveness: the MIDR Predictor achieved a maximum average Spearman’s rank correlation coefficient of $0.74$, with an average coefficient exceeding $0.91$ for severe fire scenarios. The MFSP Predictor, at the room level, attained Top-5 and Top-10 accuracies of up to $74.1\%$ and $83.2\%$, respectively. These results highlight the potential of the proposed framework to enhance fire safety assessment in buildings.
Despite these achievements, several limitations and opportunities for future research remain. The dataset used in this study, generated through a simplified approach, aligns with real-world scenarios but may not fully capture the complexity of structural geometry or thermal fields under fire conditions. Incorporating real-world measurement data or use of the Fire Dynamics Simulator (FDS) could improve model accuracy. Challenges related to limited data volume could be addressed through TL, domain adaptation, or data augmentation, leveraging trained models from this study as a foundation.
From a modeling perspective, the depth of GNN layers was constrained to match the number of building stories, successfully mitigating over-smoothing issues. However, this constraint may hinder communication between nodes in low-rise, large-area buildings, potentially limiting the model’s ability to capture horizontal fire spread. Future work could explore alternative architectures, e.g., dummy nodes or enhanced message-passing mechanisms, to overcome such challenges.
Additionally, exploring theoretical connections between GNN performance and FEA tools could provide a more robust theoretical foundation for applying GNNs in structural analysis. Expanding the framework to incorporate other data types, such as image data or OpenSees scripts, could further enhance predictive accuracy and applicability.
Beyond the MFSP prediction task, the proposed framework demonstrates potential for broader applications within civil engineering. The differentiable nature of the MIDR Predictor could be leveraged to tackle other structural challenges, such as identifying the weakest points under gravity and/or lateral loads or detecting regions most susceptible to corrosion. These extensions highlight the versatility of the framework and its potential to address a wide range of structural analysis problems, fostering future innovations in structural safety and resilience.

### v5

@@ VII-B2 Evaluation results
TABLE VIII: Key CDF values (%) of different metrics to evaluate the MFSP predictor (numbers in boldface indicate best results).
@@
Comparing the De novo training and TL with GNN methods, the latter outperforms the former in all metrics. For both small and large predictors, TL with GNN achieves lower average distance error, average room distance error, and average room rank, while also attaining higher average MIDR values compared to De novo training using the same loss function. This indicates that the TL with GNN approach effectively leverages the information embedded in the MIDR predictor’s GNN module to enhance the performance of the MFSP predictor.
@@
We compare the MSE loss with Hybrid loss within the TL with GNN method, using the Hybrid ($w_{1}=50$) case as an example. For the average MIDR metric, the Hybrid loss explicitly calculates MIDR, achieving a higher value than the MSE loss by up to $(2.46-2.40)\left/2.40\right.=2.50\%$ for the small predictor, with a smaller improvement for the large predictor. Conversely, as the MSE loss is computed based on the pseudo ground truth MFSP (similar to the distance error metric), it achieves a lower average distance error than the that of the Hybrid loss by up to $16\%$ for the large predictor. Interestingly, despite the average room distance error and room rank being calculated relative to the pseudo ground truth MFSP, the Hybrid loss performs similar to or even better than the MSE loss. This is likely attributed to the pseudo labels being accurate at the room-level granularity, even though the MFSP is not necessarily located at the center of the room. As a result, the MSE and Hybrid loss methods show varying performances across the distance error and room distance error metrics.
@@
The Hybrid loss introduces a trade-off between achieving higher MIDR and lower distance error by adjusting the weights $w_{1}$ and $w_{2}$, which reflect the degree of reliance on the pseudo labels. In this analysis, the weight for the MSE term is fixed at $w_{2}=1$. By varying the weight for the MIDR term ($w_{1}$) from 10 to 100, the reliance on pseudo labels decreases as $w_{1}$ increases. This trade-off is evident when comparing results for $w_{1}=10$ and $w_{1}=50$, where $w_{1}=10$ results in about $10\%$ lower distance error but also yields a slightly lower MIDR, reduced by about $1\%$. When $w_{1}$ is set to a very large value, such as $w_{1}=100$, the reliance on the pseudo labels becomes minimal, and the difficulty of training an “argmaxer”, as discussed in Section V-A, negatively impacts the MFSP predictor’s performance. For the average MIDR metric, the performance with $w_{1}=100$ is worse than that achieved with $w_{1}=50$. Although the average MIDR metric does not vary monotonically, it is noteworthy that it changes only slightly with $w_{1}$, indicating that $w_{1}$ is a robust choice.
@@
Taking TL with GNN and Hybrid loss ($w_{1}=50$) as an example, the average room distance error reaches 1.15, i.e., the predicted MFSP is, on average, located within 1.15 rooms of the ground truth MFSP. Additionally, the average room rank is $6.73$ indicating that the predicted MFSP is approximately the 6th or 7th most fire-sensitive room among all rooms in the building, with each structure containing an average of over 87 rooms herein. These low values suggest that the MFSP predictor accurately identifies the MFSPs in the building.
Figure 11: Illustrations of positions of ground truth and predicted MFSPs in three views of the building of ex2.
Quantile performance
@@
To further illustrate the performance of the MFSP predictor, we plotted the Cumulative Distribution Function (CDF), defined as ${\cal{F}}_{X}(a)={\cal{P}}(X\leq a)$, where $\cal{P}$ indicates the probability and $X$ is a random variable representing the different metrics, i.e., distance error, room distance error, and room rank. Figures 10(a to c) display the results for the small predictor, while Figures 10(d to f) present the results for the large predictor. The key CDF values are summarized in Table VIII. In Figure 10, the CDF curves of TL with GNN generally lie above those of De novo training, demonstrating that the former outperforms the latter in most cases. Besides, TL with GNN requires much less computational resources in training because it leverages a pre-trained GNN instead of training from scratch. The pre-trained model provides a well-initialized starting point, requiring only fine-tuning to adapt to the target task. This further highlights the efficiency and superiority of the approach of TL with GNN.
@@
From Table VIII, for the small network using TL with GNN and Hybird loss ($w_{1}=50$), $65.40\%$ of the cases achieve a distance error of less than $5$ meters. When the threshold is extended to $10$ meters, this ratio increases to $83.21\%$, surpassing the $82.19\%$ achieved with De novo training. Regarding room distance error, $\tilde{e}_{m}$, the large network using TL with GNN and Hybird loss achieves the highest probability of $73.93\%$ for $\tilde{e}_{m}\leq\sqrt{2}$, indicating that the predicted MFSP is either within the same room as the pseudo ground truth MFSP or in an adjacent room. Notably, for normalized rooms treated as unit-length squares (Section VII-B1), the longest distance between the centers of two diagonally adjacent rooms is $\sqrt{2}$. When the error threshold is further relaxed to $\tilde{e}_{m}\leq 2$, this probability increases to $80.32\%$, reflecting a high level of accuracy. For room rank, the large network employing TL with GNN and Hybird loss achieves the highest Top-5 accuracy, defined as the fraction of instances where the predicted MFSR ranks among the top 5 pseudo ground truth MFSRs, i.e., ${\cal{P}}\left(r_{m}\leq 5\right)$. This accuracy reaches $74.06\%$, representing an improvement of $2.49\%$ over De novo training. When using MSE loss, the TL with GNN method achieves a Top-5 accuracy of $72.28\%$, i.e., $3.7\%$ higher than that of De novo training. These results highlight the robustness of the TL with GNN approach, particularly when paired with Hybrid loss, in accurately identify the most fire-sensitive locations within buildings.
Case Study Building of ex2
We take ex2 (Figure 6(b)) as an example to demonstrate the evaluation process of the MFSP prediction. In this case, the interior space of the building is discretized into a set of uniformly distributed grid points with approximately $\leq 2$ meters spacing in each direction, resulting in a total of $1,232$ grid points. We adopt a brute-force strategy by simulating each grid point as a potential fire point using the FEA simulation tool, and recording the corresponding MIDR for each grid point. Among all grid points, the one with the highest MIDR is considered the 3D vector of ground truth MFSP, denoted as $\text{MFSP}_{\text{gt}}=[15.97,0.45,25.53]$ meters. This requires roughly $1,232\times 48.34\,\text{s}\approx 16.54$ hr to obtain this $\text{MFSP}_{\text{gt}}$ (refer to Table VI). We then employ our proposed large MFSP predictor (TL with GNN) to predict the MFSP. Under the MSE-only training scheme, the predicted 3D vector of MFSP is $\text{MFSP}_{\text{mse}}=[13.73,3.38,25.78]$, resulting in a prediction error of $3.7$ meters. Under the Hybrid Loss ($w_{1}=50$), the predicted 3D vector of MFSP is $\text{MFSP}_{\text{hybrid}}=[16.29,0.01,24.95]$, with a much lower prediction error of $0.8$ meters. Since the MFSP predictors share the same architecture as the MIDR predictors, the inference time is still only several ms (refer to Table VI).
Figure 11 provides a visual comparison in three orthogonal planes ($X-Y$, $Y-Z$, and $X-Z$) containing $\text{MFSP}_{\text{gt}}$. In each subplot, the MIDR values corresponding to all grid points are plotted as heatmaps, along with the projections of the predicted MFSPs. Notably, the grid point closest to the $\text{MFSP}_{\text{hybrid}}$ is exactly the same point as the $\text{MFSP}_{\text{gt}}$, verifying the accuracy of the Hybrid loss method. From Figure 11(a), we observe that the MIDR distribution in the $X-Y$ plane is clearly asymmetric. This is a deliberate consequence of introducing asymmetry into the simulation setup, in order to reflect the irregularity of real-world building geometries and uncertainty in the design procedures and building operations. Figures 11(b) and (c) further illustrate the strong correlation between MIDR and the vertical position (floor level) of the fire point. The simulation results indicate that fires originating on lower floors tend to have limited impact on the overall structural stability in the relatively short simulation time. In contrast, fires occurring on higher floors are more likely to cause larger MIDRs in upper stories, posing a greater threat to the structural integrity. This insight suggests that structural strengthening for fire safety may need to be more focused on higher floors. Moreover, we note that the MIDR values do not exhibit a monotonic relationship with height. The MFSP does not occur at the highest elevation grid point, and this non-trivial characteristic is successfully captured by the MFSP predictor. The Hybrid loss method (indicated by the red star) achieves near-exact alignment with the ground truth MFSP, further validating its effectiveness.
@@
In summary, the TL with GNN approach consistently outperforms De novo training across all evaluation metrics, highlighting the effectiveness of leveraging the MIDR predictor’s GNN module to enhance the performance of the MFSP predictor. Two notable trade-offs arise when selecting the loss function. The first is between $w_{1}$ and $w_{2}$ within the Hybrid loss to include reliance on the room-level granularity of the pseudo ground truth MFSP. The second is between MSE loss and Hybrid loss for higher efficiency and accuracy. While the Hybrid loss achieves slightly better results in most metrics (except for distance error), it incurs a higher computational cost due to the additional GNN forward pass required to calculate the MIDR loss and the backward pass through the MIDR predictor to calculate the gradient of the loss $L_{\text{MIDR}}$. In contrast, the MSE loss avoids this additional computational cost, making it more efficient but potentially less accurate in certain scenarios. This trade-off between computational cost and accuracy should be carefully considered based on the specific requirements and constraints of the application. If computational efficiency is a priority, the MSE loss may be more suitable. However, if higher accuracy in identifying sensitive fire points is critical, the Hybrid loss could provide a more robust solution.
VIII Limitations & Future Work
VIII-A Performance Improvement
The proposed framework demonstrates promising results, yet there are opportunities to enhance its predictive capabilities. From a modeling perspective, the constraint of depth of GNN layers to match the number of building stories effectively mitigates over-smoothing but may limit the ability to model horizontal fire spreading in low-rise, large-area buildings. Future work could explore alternative architectures, e.g., introducing dummy nodes or improving message-passing mechanisms, to enhance cross-node communication in such scenarios. Additionally, the current implementation uses shallow MLPs for key components of the message-passing network, which may restrict the model’s ability to capture complex interaction patterns. Exploring advanced GNN variants, e.g., GraphSAGE [17] or Graph Attention Networks [40], could improve prediction accuracy by better handling multi-scale spatial dependencies. A deeper analysis of the trade-off between MSE loss and MIDR loss could also further refine the model’s performance.
Beyond a purely data-driven approach, integrating Physics-Informed Neural Networks (PINNs) could leverage physical laws to refine predictions. While PINNs have shown great promise in modeling physical systems, their effectiveness depends on well-formulated Partial Differential Equations (PDEs) —a challenge in multi-physics problems like fire scenarios. In our case, the problem involves not only generating spatial thermal loads but also solving for the building’s structural response under thermal loads. These complexities make the PDE formulation challenging, limiting the immediate application of PINNs. However, as hybrid methods and improved PDE modeling techniques evolve, incorporating PINNs into fire modeling remains a valuable research direction. Hybrid approaches that combine data-driven and physics-based methods could provide a more balanced and robust solution.
Investigating the theoretical connections between GNN performance and FEA tools could strengthen the foundation for applying GNNs in structural analysis. The proposed framework is also compatible with other advanced ML techniques. For instance, ensemble learning [1] could integrate multimodal data, such as images or $\chi$ara scripts of the building structure, and self-supervised learning —proven effective for sequential data like electroencephalograms [36, 35]— could offer insights into temporal thermal analysis. Furthermore, general techniques such as neural architecture search and hyperparameter optimization could be incorporated to enhance both accuracy and practicality of the framework.
VIII-B Dataset Generation
The lack of comprehensive datasets is a significant challenge in applying ML methods to fire-related research. Our work introduces a simplified rule-based approach for efficient fire dataset generation, but it may not fully capture the complexity of real-world structural geometries or thermal fields. Incorporating experimental measurements or advanced tools such as the FDS could improve fidelity.
Balancing fidelity, efficiency, and cost remains a key consideration. The proposed rule-based method could be refined to accommodate different structural configurations or fire patterns. For example, beyond tuning the parameters, as discussed in Section VI-D, the ISO 834 curve could be replaced with a parametric time-temperature fire curve that accounts for additional factors such as ventilation conditions and fuel load density. In terms of modeling using FEA tools, some settings are chosen for simplicity of both modeling and simulation. The squared section, limited mid-element nodes, and introduced randomness of properties used in this work ensure the generation of a large dataset with acceptable cost to train the proposed predictors and test the workflow as a proof-of-concept. Such simplifications may introduce impractical scenarios in the generated dataset. Therefore, in the case when fidelity and accuracy involving practical cases are highly desired as in final designs (instead of the considered conceptual ones herein), detailed modeling and simulation with higher demand for computational resources can be applied. However, the the concept of MFSP and the workflow of its prediction are still applicable.
While physics-based simulations like OpenSees [29, 24] facilitates large-scale training data generation, their accuracy depends on assumptions about extreme load, e.g., fire, evolution and structural responses. Rigorous experimental validation and cross-comparison studies are essential to verify simulation fidelity, particularly for complex events like fires. Future research should prioritize such validation efforts to enhance confidence in simulation outputs and downstream predictive models. Additionally, addressing data scarcity through techniques such as TL, domain adaptation, or synthetic data augmentation could further improve model robustness. Leveraging pretrained models from this study as a foundation could enable adaptation to diverse structural configurations and fire scenarios.
VIII-C Beyond Structural Analysis
The framework’s differentiable agent and modular design offer transformative potential beyond fire scenario predictions, unlocking new applications across interdisciplinary domains. A key advantage of the differentiable agent is its ability to support training auxiliary predictors (e.g., “argmaxers”) while also providing direct access to gradients that quantify how structural elements or environmental factors influence system-level responses. A natural extension of this work is structural analysis with combined loads, not only gravity and fire scenarios —for instance, assessing the impact of both earthquakes and the subsequent fire-induced structural degradation. Beyond structural analysis, the framework’s ability to identify worst-case scenarios can be generalized to other domains. For example, training a differentiable agent to simulate molecular dispersion (e.g., airborne viruses or chemical leaks) could identify the “Most Diffusion-Sensitive Point” (MDSP) —the location where contaminants spread most rapidly. Such insight could inform critical design decisions, such as ensuring that chemical laboratories or quarantine rooms are positioned away from these MDSPs, or optimizing the ventilation systems to minimize cross-contamination risks.
The agent’s gradient information can directly map element-level properties (e.g., material stiffness, airflow resistance) to global performance metrics (e.g., structural drift, contaminant concentration). This enables physics-aware sensitivity analysis, where engineers can prioritize elements with the highest gradient magnitudes —those that have the greatest impact on the system safety— when making decisions about retrofitting, material selection, or structural requirements.
By addressing these limitations and exploring the outlined directions, future work can refine the proposed framework’s accuracy, expand its scope, and unlock new possibilities for data-driven structural engineering.
IX Conclusions
@@
This study introduces a neural network-based framework for predicting the Most Fire-Sensitive Point (MFSP) in building structures, advancing fire safety assessment through a novel integration of data-driven and simulation-based approaches. The proposed framework leverages a differentiable agent model to efficiently simulate the outcome of the Finite Element Analysis (FEA) and effectively predict the Maximum Interstory Drift Ratio (MIDR) under fire conditions. This agent enables the rapid generation of MFSP-labeled data, directly supporting the training of an MFSP predictor. A rule-based temperature generation method was developed to facilitate the development of large-scale structural and fire simulation datasets and eventually forming a robust pipeline for data generation and neural network training. Key innovations of this work include:
  1. 1.
    The application of Graph Neural Networks (GNNs) to process structural data in fire scenarios.
  2. 2.
    The integration of Transfer Learning (TL) to improve prediction performance.
  3. 3.
    The introduction of an Edge Update (EU) mechanism to account for property changes in beams and columns during fire exposure.
The evaluation of the proposed framework demonstrates its effectiveness with the following two key points:
  1. 1.
    The MIDR predictor achieved a maximum average Spearman’s rank correlation coefficient of $0.74$, with an average correlation exceeding $0.91$ for severe fire scenarios.
  2. 2.
    The MFSP predictor, at the room level, attained Top-5 and Top-10 accuracies of up to $74.1\%$ and $83.2\%$, respectively.
These results highlight the proposed framework potential to enhance building fire safety assessment by efficiently identifying critical structural vulnerabilities under fire conditions, particularly during the preliminary stage of the building design.